# Normality Calibration in Semi-supervised Graph Anomaly Detection

### v1

@@ 5.3 Ablation Study
In this section, we perform an ablation study to evaluate the contribution of each component in GraphNC. To this end, several variants of GraphNC are introduced. (1) Only Teacher (OT) directly uses the pre-trained semi-supervised teacher model to derive the anomaly score. (2) OT+NormReg incorporates NormReg into the pre-training of the teacher model. (3) OT+NormReg+Finetune leverages NormReg to fine-tune the pre-trained teacher. (4) OT+ScoreDA includes the student model and adds ScoreDA on top of OT (i.e., optimizing GraphNC using ScoreDA only). (5) OT+ScoreDA+NormReg* applies NormReg on top of OT+ScoreDA but it applies the NormReg loss on all nodes instead of the labeled normal nodes only. OT+ScoreDA+NormReg is the default full model using both ScoreDA and NormReg. The experimental results are shown in Table 3. We observe that OT+ScoreDA enhances the performance of OT, achieving the second-best results on four datasets in terms of AUPRC. This demonstrates that the ScoreDA component can effectively calibrate the normality learned from the teacher, ensuring better separation between normal and abnormal classes. However, OT+ScoreDA underperforms the full model (the last row) on all datasets, highlighting that the negative impact of inaccurate anomaly scores from the teacher model can be effectively mitigated by the NormReg component. On the other hand, applying NormReg directly to the teacher via either joint optimization (OT+NormReg) or finetuning (OT+NormReg+Finetune) can improve OT to some extent. This justifies the effectiveness of NormReg from a different perspective, but NormReg works best when working in the student model and combining with ScoreDA due to its joint effects in minimizing the detection errors discussed in our theoretical analysis. Besides, we also conduct the ablation study on other teacher models to further analyze the effectiveness of each module in GraphNC. The results can be found in the App. C.2
5.4 Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of GraphNC w.r.t. the loss moderator $\alpha$, mask ratio $\omega$, and training size $R$. A detailed analysis on time complexity can be found App. E.
Performance w.r.t. the training size $R$. To explore the impact of training size $R$, we compare GraphNC with four semi-supervised GAD methods using varying numbers of training normal nodes, as shown in Figs.4 (e-f). Similar with other competing methods, with the increase of training size $R$, our model GraphNC generally performs better and it maintains the superiority over other competing methods, demonstrating the effectiveness of GraphNC using different scales of labeled training data.
6 Conclusion
In this paper, we propose GraphNC, a generic graph normality calibration framework designed to enhance existing semi-supervised GAD methods, achieving substantial and consistent performance improvements across different types of existing approaches. GraphNC is composed of two key components: ScoreDA and NormReg. The ScoreDA module aligns the anomaly scores with the anomaly score distribution produced by the teacher model, effectively pushing the anomaly scores for the normal and abnormal classes toward opposite ends. This alignment ensures a clearer separation between normal and abnormal nodes in the score space, thereby leading to a lower FPR and FNR. The NormReg module is introduced to reduce the impact of inaccurate scores in the anomaly score distribution served by the teacher, enabling GraphNC to learn more compact normal representations. GraphNC is comprehensively verified by empirical results on six GAD datasets, which show that GraphNC consistently enhances performance of three different types of teacher models, and it obtains better performance if the teacher model is stronger.
Limitation and Future work. A potential limitation of GraphNC is that it relies on an assumption that the pre-trained teacher model performs fairly well to provide informative supervision signals for the ScoreDA component. Future work will explore other sources, such as some labeled anomaly data or expert knowledge, as supplementary supervision to reduce its reliance on the teacher model.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. The datasets used in our experiments are all publicly available, and the construction and analysis of the data are presented in detail in App. A. The implementation details, including model configurations and training hyperparameters, are thoroughly documented in Sec. 5. The source code has been submitted as supplementary materials.
Appendix A Detailed Description of Datasets
  • YelpChi (Dou et al., 2020): Similar to Amazon-all, YelpChi-all incorporates three types of edges: R-U-R (reviews written by the same user), R-S-R (reviews of the same product with identical star ratings), and R-T-R (reviews of the same product posted within the same month). YelpChi-all is then constructed by merging these different relations into a single unified relation (Chen et al., 2022a; Qiao & Pang, 2023).
Appendix B Description of Baselines
  • GAAN (Chen et al., 2020): It leverages a generative adversarial network where fake graph nodes are generated. Then the covariance matrix for real nodes and fake nodes are computer to enhace the node. Finally, a discriminator is trained to recognize whether two connected nodes are from a real or fake node.
  • TAM (Qiao & Pang, 2023): It first reveals the ’one class homophily’ and introduces the new anomaly measure score ’local node affinity’. As the local node affinity can not be directly obtained based on the embedding, it is optimized on truncated graphs where non-homophily edges are removed iteratively. The local node affinity is calculated on the learned representations on the truncated graphs.
Appendix C Additional Experimental Results
Table 4: Key statistics of GAD datasets.
Datasets Amazon T-Finance Reddit YelpChi Tolokers Photo
#Nodes 11,944 39,357 10,984 45,941 11,758 7,484
#Edges 4,398,392 21,222,543 168,016 3,846,979 519,000 119,043
#Attributes 25 10 64 32 10 745
Anomaly 9.5% 4.6% 3.3% 14.52% 21.8% 4.9%
C.1 Sore distribution visualization on other teacher model
In this section, we visualized the score distributions of the reconstruction-based method DOMINANT and the one-class classification method OCGNN, along with their corresponding GraphNC-enabled models on the Amazon. As shown in the Fig. 6, we observe that the GraphNC-enabled model achieves better separation between normal and abnormal classes, further demonstrating that GraphNC can substantially and consistently enhance performance.
(a) DOMINANT
(b) GraphNC(DOMINANT)
Figure 5: The score distribution of DOMINANT along the corresponding NomrDR enabled DOMINANT on Amazon (Dou et al., 2020)
(a) OCGNN
(b) GraphNC(OCGNN)
Figure 6: The score distribution of OCGNN along the corresponding GraphNC enabled OCGNN on Amazon (Dou et al., 2020)
C.2 Ablation study on other teacher models
@@
We also conduct the ablation study on the GraphNC-enabled other two teacher models to demonstrate the effectiveness of each module in the GraphNC. We employ the reconstruction-based method, DOMINANT, and one-class classification-based method, OCGNN as the teacher model and make a comparison with the variant of GraphNC. The experimental results are shown in the Table. 5 We observe that GraphNC outperforms the corresponding variant on most of the datasets for both DOMINANT and OCGNN. This further demonstrates the flexibility and effectiveness of GraphNC in enhancing existing semi-supervised methods.
@@
Table 5: AUROC and AUPRC results comparison of the proposed method GraphNC and its variants.
@@ Appendix D Theoretical Analysis
Appendix E Time Complexity Analysis
@@ Appendix F Algorithm
The algorithm of GraphNC is summarized in Algorithm 1
@@
Input: Graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$; Pre-trained Semi-supervised GAD Model $\mathcal{F}_{\mathcal{T}}$; Student Model $\mathcal{F}_{S}$ with parameters $\Phi$; Labeled Normal nodes $\mathcal{V}_{l}\subset\mathcal{V}$; Number of training epochs $E$; Trade-off weight $\alpha$.
Output: Anomaly score $y^{\mathcal{S}}_{i}$ for each node $v_{i}\in\mathcal{V}$.
1: Initialize parameters $\Theta$ for the student model $\mathcal{F}_{S}$.
2: Obtain the anomaly scores distribution $\mathcal{Y}^{\mathcal{T}}=\mathcal{F}_{\mathcal{T}}{(\mathbf{X},\mathcal{G};\Theta)}=\{y_{1}^{\mathcal{T}},y_{2}^{\mathcal{T}},...,y_{N}^{\mathcal{T}}\}$ from the teacher model.
3: Create an augmented feature matrix $\mathbf{X}^{\prime}$ by applying $\operatorname{RandomMask}$ to features of nodes in $\mathcal{V}_{l}$.
4:for $e=1$ to $E$ do
5:  for $v_{i}$ in $\mathcal{V}$ do
6:   Obtain the representation for node $v_{i}$: $\mathbf{h}_{i}^{\mathcal{S}}\leftarrow\mathcal{F}_{GNN}(\mathbf{x}_{i},\mathcal{G},\mathbf{X};\Omega)$.
7:   Obtain the augmented representation for node $v_{i}$: $\tilde{\mathbf{h}_{i}^{\mathcal{S}}}\leftarrow\mathcal{F}_{GNN}(\tilde{\mathbf{x}_{i}},\tilde{\mathcal{G}},\tilde{\mathbf{X}};\Omega)$.
8:   Obtain the anomaly score of student model for node $v_{i}$: $y^{\mathcal{S}}_{i}=\mathcal{F}_{MLP}({\bf{h}}_{i}^{\mathcal{S}};\phi)$.
9:  end for
10:  // — Anomaly Score Distribution Alignment (ScoreDA) — $\mathcal{L}_{ScoreDA}=\frac{1}{|{{\cal V}}|}\sum_{v_{i}\in{\mathcal{V}}}\|y^{\mathcal{S}}_{i}-y^{\mathcal{T}}_{i}\|_{2}^{2}$.
11:  // — Perturbation-based Regularization (NormaReg) —
12:  ${{\cal L}_{{NormReg}}}=-\frac{1}{{\left|{{{\cal V}_{l}}}\right|}}\sum\limits_{{v_{i}}\in{\cal V}_{l}}{\left\|{{{\bf{h}}_{i}^{\mathcal{S}}}-{{\widetilde{\bf{h}}}_{i}^{\mathcal{S}}}}\right\|_{2}^{2}}$.
13:  // — Total Loss Function —
14:  $\mathcal{L}_{GraphNC}\leftarrow\mathcal{L}_{ScoreDA}+\alpha\cdot\mathcal{L}_{NormReg}$.
15:  Minizing $\mathcal{L}_{GraphNC}$ and update parameters of student model, $\Phi=\{\Omega,\phi\}$ by gradient descending .
16:end for
17:return Anomaly scores $S(v_{i})=\mathcal{F}_{\mathcal{S}}(v_{i},{\bf{x}}_{i},\mathcal{G},\Phi^{*}),$ for each node $v_{i}\in\mathcal{V}$

### v2

@@ C.2 Ablation study on all datasets
In this section, we perform an ablation study to evaluate the contribution of each component in GraphNC. To this end, several variants of GraphNC are introduced. (1) Only Teacher (OT) directly uses the pre-trained semi-supervised teacher model to derive the anomaly score. (2) OT+NormReg incorporates NormReg into the pre-training of the teacher model. (3) OT+NormReg+Finetune leverages NormReg to fine-tune the pre-trained teacher. (4) OT+ScoreDA includes the student model and adds ScoreDA on top of OT (i.e. , optimizing GraphNC using ScoreDA only). (5) OT+ScoreDA+NormReg* applies NormReg on top of OT+ScoreDA but it applies the NormReg loss on all nodes instead of the labeled normal nodes only. OT+ScoreDA+NormReg is the default full model using both ScoreDA and NormReg. The experimental results are shown in Table 5. We observe that OT+ScoreDA enhances the performance of OT, achieving the second-best results on four datasets in terms of AUPRC. This demonstrates that the ScoreDA component can effectively calibrate the normality learned from the teacher, ensuring better separation between normal and abnormal classes. However, OT+ScoreDA underperforms the full model (the last row) on all datasets, highlighting that the negative impact of inaccurate anomaly scores from the teacher model can be effectively mitigated by the NormReg component. On the other hand, applying NormReg directly to the teacher via either joint optimization (OT+NormReg) or finetuning (OT+NormReg+Finetune) can improve OT to some extent. This justifies the effectiveness of NormReg from a different perspective, but NormReg works best when working in the student model and combining with ScoreDA due to its joint effects in minimizing the detection errors discussed in our theoretical analysis. Besides, we also conduct the ablation study on other teacher models to further analyze the effectiveness of each module in GraphNC. The results can be found in the App. C.3
C.3 Ablation study on other teacher models
@@
We also conduct the ablation study on the GraphNC-enabled other two teacher models to demonstrate the effectiveness of each module in the GraphNC. We employ the reconstruction-based method, DOMINANT, and one-class classification-based method, OCGNN as the teacher model and make a comparison with the variant of GraphNC. The experimental results are shown in the Table 6 We observe that GraphNC outperforms the corresponding variant on most of the datasets for both DOMINANT and OCGNN. This further demonstrates the flexibility and effectiveness of GraphNC in enhancing existing semi-supervised methods.
@@
Table 6: AUROC and AUPRC results comparison of the proposed method, GraphNC and its variants.
@@
C.4 The performance of GraphNC when the teacher performs poorly
To evaluate the performance of GraphNC when the teacher performs poorly (AUROC $\leq$ 0.5), we inject anomalies into the labeled normal nodes to deliberately degrade the teacher’s performance. When the teacher model performs poorly, we still apply score distribution alignment and NormReg to achieve normality calibration. The experimental results using GGAD as teacher model are shown in Table 7.
From the results, we observe that when the teacher’s performance is poor, e.g., the AUROC is lower than 0.5 on some datasets, GraphNC can largely enhance the performance. The main reason is that the errors introduced by teacher distillation mainly affect a small portion of nodes, for example, some normal nodes may receive slightly higher scores (while most normal nodes still remain low), and some anomalies may obtain lower scores. However, NormReg relies on the explicitly labeled normal nodes, rather than the teacher’s predictions, to enforce the compactness of the normality cluster. Consequently, during the ScoreDA process, the normal cluster continues to be tightened, while the correctly predicted normal nodes guide the adjustment of other normal nodes. This helps correct some of the teacher’s mistakes, ensuring that GraphNC can still achieve performance gains even when the teacher performs poorly.
Moreover, as shown in Table 2 of the paper, we observe that DOMINANT and OCGNN perform poorly on certain datasets. For instance, on Photo, DOMINANT exhibits notably low performance, suggesting that the teacher’s label distribution becomes unreliable during score alignment. GraphNC leverages NormReg to prevent overfitting to these inaccurate supervisory signals and to better exploit the limited yet accurate normal supervision, thereby improving overall performance. Finally, we summarize the average improvement of GraphNC over the teacher model across different teacher score ranges, as presented in the Table 8.
Table 7: AUROC and AUPRC results of GraphNC and GGAD* (GGAD is degraded by injecting anomalies into the labeled normal set).
Metric Method Dataset
Amazon T-Finance Reddit YelpChi Tolokers Photo Avg.
AUROC GGAD* 0.6336 0.5132 0.4483 0.5130 0.4504 0.4465 0.5008
GraphNC 0.7769 0.7450 0.5303 0.5685 0.6291 0.5970 0.6411
AUPRC GGAD* 0.1171 0.0443 0.0276 0.1603 0.1946 0.0856 0.1049
GraphNC 0.2274 0.0930 0.0375 0.1823 0.2930 0.1145 0.1580
C.5 Analysis on true anomalous instances recognition
To offer more insights into the dependency between our model and the teacher model, we provide statistics for the number of anomalous instances which are accurately detected by the teacher (GGAD) versus those detected by GraphNC in Fig. 7. We can observe that the number of anomalous instances accurately detected by GraphNC is consistently and substantially higher than the teacher model across different datasets where the teacher model’s performance varies largely. This indicates that GraphNC can go far beyond the supervision from the teacher model to detect anomalies that the teacher model missed. We attribute this capability to the NormReg component and its synergy with the ScoreDA component.
Figure 7: The number of true anomalous instances that are accurately detected by the teacher (GGAD) versus those detected by GraphNC.
Table 8: Statistics of the average AUROC improvement of GraphNC w.r.t. different accuracy levels of supervision from the teacher model GGAD.
[…]
GraphNC 0.1345 0.1128 0.0764 0.0204
C.6 Statistics of the average AUROC improvement of GraphNC w.r.t. different accuracy levels
[…]
@@ Appendix D Theoretical Analysis
Appendix E Efficiency Analysis
E.1 Time Complexity Analysis
@@
E.2 Runtime Analysis
We provide a runtime comparison between the standalone teacher model and the model augmented with GraphNC below. The training and inference time comparison are shown in Fig. 8. It is clear that GraphNC is highly efficient for training, as its training process builds directly upon the pretrained teacher model. In addition, GGAD and GraphNC are both very efficient for inference.
(a) Running times comparison of training
(b) Running times comparison of inference
Figure 8: Training and inference runtimes (in seconds) on the six datasets for GGAD and GraphNC.
@@ Appendix F Algorithm
The algorithm of GraphNC is summarized in Algorithm 1.
@@
1:Input: Graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$; pre-trained semi-supervised GAD model $\mathcal{F}_{\mathcal{T}}$; student model $\mathcal{F}_{S}$ with parameters $\Phi$; labeled normal nodes $\mathcal{V}_{l}\subset\mathcal{V}$; training epochs $E$; weight $\alpha$.
2:Output: Anomaly scores $\{y_{i}^{\mathcal{S}}\}_{v_{i}\in\mathcal{V}}$.
3: Initialize parameters $\Phi$ for the student model $\mathcal{F}_{S}$.
4: Obtain teacher score distribution $\mathcal{Y}^{\mathcal{T}}=\mathcal{F}_{\mathcal{T}}(\mathbf{X},\mathcal{G})=\{y_{1}^{\mathcal{T}},\dots,y_{N}^{\mathcal{T}}\}$.
5: Create augmented features $\tilde{\mathbf{X}}$ by applying $\operatorname{RandomMask}$ to features of nodes in $\mathcal{V}_{l}$.
6:// GraphNC Training
7:for $e=1$ to $E$ do
8:  Compute student representations $\mathbf{H}^{\mathcal{S}}\leftarrow\mathcal{F}_{GNN}(\mathbf{X},\mathcal{G};\Omega)$.
9:  Compute augmented representations $\tilde{\mathbf{H}}^{\mathcal{S}}\leftarrow\mathcal{F}_{GNN}(\tilde{\mathbf{X}},\mathcal{G};\Omega)$.
10:  Compute student scores $\mathcal{Y}^{\mathcal{S}}\leftarrow\mathcal{F}_{MLP}(\mathbf{H}^{\mathcal{S}};\phi)$.
11:  // Anomaly Score Distribution Alignment (ScoreDA)
12:  $\mathcal{L}_{ScoreDA}\leftarrow\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\|y_{i}^{\mathcal{S}}-y_{i}^{\mathcal{T}}\|_{2}^{2}$.
13:  // Perturbation-based Normality Regularization (NormReg)
14:  $\mathcal{L}_{NormReg}\leftarrow\frac{1}{|\mathcal{V}_{l}|}\sum_{v_{i}\in\mathcal{V}_{l}}\|\mathbf{h}_{i}^{\mathcal{S}}-\tilde{\mathbf{h}}_{i}^{\mathcal{S}}\|_{2}^{2}$.
15:  $\mathcal{L}_{GraphNC}\leftarrow\mathcal{L}_{ScoreDA}+\alpha\cdot\mathcal{L}_{NormReg}$.
16:  Update $\Phi=\{\Omega,\phi\}$ by minimizing $\mathcal{L}_{GraphNC}$.
17:end for
18:// GraphNC Anomaly Score Inference
19:for $i=1$ to NN do
20:  Load the test graph $\mathcal{G}$ and optimal parameters $\Phi^{*}$ for the student model $\mathcal{F}_{S}$.
21:  Anomaly Scoring $y_{i}^{\mathcal{S}}=\mathcal{F}_{\mathcal{S}}(v_{i},{\bf{x}}_{i},\mathcal{G},\Phi^{*})$
22:end for
23:return $\{y_{i}^{\mathcal{S}}\}_{v_{i}\in\mathcal{V}}$.